# SDXS: Real-Time One-Step Latent Diffusion Models with Image Conditions

### v1

@@ 1 Introduction
Table 1: Latency Comparison between SD v2.1 base and our proposed efficient diffusion models on generating $512\times 512$ images with batch size = 1. SD v2.1 base Text Encoder U-Net Image Decoder #Parameters 0.33B 0.87 B 50 M Latency (ms) 1.1 16 18 NFEs 2 32 1 Total (ms) 2.2 512 18 SDXS-512 Text Encoder U-Net Image Decoder #Parameters 0.33B 0.32 B 1.2 M Latency (ms) 1.1 6.3 1.7 NFEs 1 1 1 Total (ms) 1.1 6.3 1.7
Table 2: Latency Comparison between SDXL and our proposed efficient diffusion models on generating $1024\times 1024$ images with batch size = 1. SDXL Text Encoder U-Net Image Decoder #Parameters 0.80B 2.56 B 50 M Latency (ms) 1.8 56 73 NFEs 2 32 1 Total (ms) 3.6 1792 73 SDXS-1024 Text Encoder U-Net Image Decoder #Parameters 0.80B 0.74 B 1.2 M Latency (ms) 1.8 24 6.1 NFEs 1 1 1 Total (ms) 1.8 24 6.1
In this paper, we provide a more comprehensive exploration of the aforementioned challenges. Initially, we focus on reducing the size of the VAE [35] decoder and U-Net [36], both are resource-intensive components in DM sampling. We train an extremely lightweight image decoder to mimic the original VAE decoder’s output through a combination of output distillation loss and GAN loss. Following this, we leverage the block removal distillation strategy [16] to efficiently transfer the knowledge from the original U-Net to a more compact version, effectively removing the majority of modules that contribute to latency. To reduce the NFEs, we propose a fast and stable training method. First, we suggest straightening the sampling trajectory and quickly finetuning the multi-step model into a one-step model by replacing the distillation loss function with the proposed feature matching loss. Then, we extend the Diff-Instruct training strategy [37], using the gradient of the proposed feature matching loss to replace the gradient provided by score distillation in the latter half of the timestep. We name our proposed method SDXS to admire SDXL [2], and provide two versions: $512\times 512$ and $1024\times 1024$, developed based on SD v2.1 base and SDXL, respectively. As shown in Tables 1 and 2, SDXS demonstrates efficiency far surpassing that of the base models, even achieving image generation at 100 FPS for $512\times 512$ images and 30 FPS for $1024\times 1024$ images on the GPU. Finally, to ease the application of our optimized model to tasks involving image-conditioned generation, we have adapted the block removal distillation strategy for ControlNet [5]. Then, we extend our proposed training strategy to the training of ControlNet, relying on adding the pretrained ControlNet to the score function.
@@ 2.2 Diff-Instruct
where $\bm{x}_{0}=g_{\theta}(\bm{z})$ denotes the sample generated by the one-step generator being trained to accept a randomly initialized latent $z$, and $\bm{s}_{\phi}$ and $\bm{s}_{p_{t}}$ denote the score functions of the DM trained online on the generated data and the pretrained DM, respectively. Diff-Instruct uses this gradient directly to update the generator, and when the outputs of the two score functions agree, the marginal distribution of the one-step generator output is consistent with the marginal distribution of the pretrained DM.
@@ 3.1 Architecture Optimizations
where $\lambda_{F}$ balances the two loss terms. Different from BK-SDM, we exclude the original denoising loss. Our models are miniaturized based on SD-2.1 base (resolution $512\times 512$) and SDXL-1.0 base (resolution $1024\times 1024$). For SD-2.1 base, we remove the middle stage, the last stage of the downsampling stages and the first stage of the upsampling stages, and remove the Transformer blocks of the highest resolution stages. For SDXL-1.0 base, we remove most of the Transformer blocks.
@@ 3.2 One-Step Training
A straightforward approach to training a one-step model involves initializing noises $\bm{\epsilon}$ and employing an Ordinary Differential Equation (ODE) sampler $\psi$ to sample and obtain the generated images $\bm{\hat{x}}_{0}$, thereby constructing noise-image pairs. These pairs then serve as inputs and ground truth for the student model during training. This method, however, often results in the production of blurry and low-quality images. The underlying issue is the crossings in the sampling trajectories of noise-image pairs generated using the ODE sampler from a pretrained DM, leading to an ill-posed problem. Rectified Flow [29, 30] tackles this challenge by straightening the sampling trajectories. It replaces the training objective and proposes a ’reflow’ strategy to refine the pairing, thus minimizing trajectory crossings. Yet, given Rectified Flow’s complexity and lengthy training time, an efficient method to refine a multi-step model into a one-step version is desirable. We note that the crossing of sampling trajectories can lead to a one noise input corresponding to multiple ground-truth images, causing the trained model generating an image that is a weighted sum of multiple feasible outputs with weights $w(y)$:
@@
Segmented Diff-Instruct.
@@
$\lambda_{FM}$ is used to balance gradients of the two segments, and $\alpha\in[0,1]$. we intentionally set $\alpha$ close to 1 and $\lambda_{FM}$ at a high value to ensure the model’s output distribution smoothly aligns the predicted distribution by the pretrained score function. After achieving significant overlap in the probability densities of the model’s output and the target distributions, we gradually lower both $\alpha$ and $\lambda_{FM}$. This gradual reduction finetunes the model’s focus, enhancing its alignment with the target distribution for more detailed matching. Figure 3 visually depicts our training strategy, where the offline DM represents the U-Net of a pre-trained DM like SDXL, and the online DM is initialized from the offline DM and finetuned on the generated images through Eq. (1). In practice, the online DM and student DM are trained alternately. In the mid-term of training, we switch the teacher DM from the pretrained model to the trained miniaturized one-step model itself, further reducing training overhead.
@@
Our approach can also be adapted for training ControlNet, enabling the tiny one-step model to incorporate image conditions into its image generation process, as depicted in Figure 4. Compared to the base model for text-to-image generation, the model trained here is the distilled ControlNet that accompanies the tiny U-Net mentioned earlier, and the parameters of the U-Net are fixed during training. Importantly, we need to extract the control images from the images sampled by the teacher model, rather than from the dataset images, to ensure that the noise, target image, and control image form a pairing triplet. Furthermore, the original multi-step U-Net’s accompanying pretrained ControlNet is integrated with both the online U-Net and offline U-Net but does not participate in training. Similar to the text encoder, the function is confined to serving as a pretrained feature extractor. The core idea of this method is that the student DM cannot perfectly mimic the sampling results of the teacher DM due to its smaller model capacity and one-step sampling. Therefore, to further reduce $\lambda_{FM}$, the trained ControlNet is to learn to utilize the control images extracted from the target images. At the same time, the score distillation part encourages the model to match the marginal distributions, enhancing the contextual relevance of the generated images. Notably, we found that replacing a portion of the noise input to U-Net with freshly reinitialized noise can enhance control capabilities.
@@ 4 Experiment
SD v1.5 [1] $512\times 512$ 860 M DPM-Solver++ 16 276 24.28 31.84
SD v1.5-LCM [28] $512\times 512$ 860 M LCM 4 84 34.74 30.85
SD Turbo [32] $512\times 512$ 865 M - 1 35 26.50 33.07
Tiny SD [16] $512\times 512$ 324 M DPM-Solver++ 16 146 31.16 31.06
SDXS-512 (ours) $512\times 512$ 319 M - 1 9 28.21 32.81
@@
SDXL [2] $1024\times 1024$ 2.56B Euler 32 1869 24.60 33.77
SDXL-LCM [28] $1024\times 1024$ 2.56B LCM 4 299 28.87 32.45
SDXL Lightning [33] $1024\times 1024$ 2.56B - 1 131 29.26 32.09
Vega [17] $1024\times 1024$ 0.74B Euler 32 845 29.24 32.99
SDXS-1024 (ours) $1024\times 1024$ 0.74B - 1 32 30.92 32.32
@@
Implementation Details. Our code is developed based on diffusers library [^1] https://github.com/huggingface/diffusers. Because we cannot access the training datasets of SD v2.1 base and SDXL, the entire training process is almost data-free, relying solely on prompts available from publicly accessible dataset [57]. When necessary, we use open-source pretrained models in conjunction with these prompts to generate corresponding images. To train our model, we configure the training mini-batch size to range from 1,024 to 2,048. To accommodate this batch size on the available hardware, we strategically implement gradient accumulation when necessary. It’s important to note that we found that the proposed training strategy results in models generating images with less texture. Therefore, after training, we utilize GAN loss in conjunction with extremely low-rank LoRA [34] for a short period of fine-tuning. When GAN loss is needed, we use Projected GAN loss from StyleGAN-T [58], and the basic settings are consistent with ADD [32]. For the training of SDXS-1024, we use Vega [17], a compact version of SDXL, as the initialization for both online DM and offline DM to reduce training overhead.
@@ 4.1 Text-to-Image Generation
We report the quantitative results, i.e., FID [59] and CLIP scores [60], on MS-COCO 2017 validation set [61] for evaluation. Due to the strong assumption of Gaussian distribution, FID is not a good indicator for measuring image quality [2], as it is significantly affected by the diversity of generated samples, but we still report it as prior works do. Table 3 shows the performance comparison on MS-COCO 2017 5K subset. Despite a noticeable downsizing in both the sizes of the models and the number of sampling steps required, the prompt-following capability of SDXS-512 remains superior to that of SD v1.5. Moreover, when compared to Tiny SD, another model designed for efficiency, the superiority of SDXS-512 becomes even more pronounced. This observation is consistently validated in the performance of SDXS-1024 as well.
@@ 5 Conclusion
This paper explores the distillation of large-scale diffusion-based text-to-image generation models into efficient versions that enable real-time inference on GPUs. Initially, we employ knowledge distillation to compress both the U-Net architecture and the image decoder. Subsequently, we introduce a novel training strategy that leverages feature matching and score distillation to reduce the sampling process to one step. This approach allows for the real-time generation of $1024\times 1024$ images on a single GPU, maintaining quality comparable to original models. Moreover, the training methodology we propose can also adapt to tasks involving image-conditioned generation, eschewing the direct adaptation of the pretrained ControlNet. We believe that the deployment of efficient image-conditioned generation on edge devices represents a promising avenue for future research, with plans to explore additional applications such as inpainting and super-resolution.

### v2

@@ 1 Introduction
Table 1: Latency Comparison between SD v2.1 base and our proposed efficient diffusion models on generating $\times 512$ 512 images with batch size = 1.
SD v2.1 base Text Encoder U-Net Image Decoder
#Parameters 0.33B 0.87 B 50 M
Latency (ms) 1.1 16 18
NFEs 2 32 1
Total (ms) 2.2 512 18
SDXS-512 Text Encoder U-Net Image Decoder
#Parameters 0.33B 0.32 B 1.2 M
Latency (ms) 1.1 6.3 1.7
NFEs 1 1 1
Total (ms) 1.1 6.3 1.7
Table 2: Latency Comparison between SDXL and our proposed efficient diffusion models on generating $\times 1024$ 1024 images with batch size = 1.
SDXL Text Encoder U-Net Image Decoder
#Parameters 0.80B 2.56 B 50 M
Latency (ms) 1.8 56 73
NFEs 2 32 1
Total (ms) 3.6 1792 73
SDXS-1024 Text Encoder U-Net Image Decoder
#Parameters 0.80B 0.74 B 1.2 M
Latency (ms) 1.8 24 6.1
NFEs 1 1 1
Total (ms) 1.8 24 6.1
In this paper, we provide a more comprehensive exploration of the aforementioned challenges. Initially, we focus on reducing the size of the VAE [35] decoder and U-Net [36], both are resource-intensive components in DM sampling. We train an extremely lightweight image decoder to mimic the original VAE decoder’s output through a combination of output distillation loss and GAN loss. Following this, we leverage the block removal distillation strategy [16] to efficiently transfer the knowledge from the original U-Net to a more compact version, effectively removing the majority of modules that contribute to latency. To reduce the NFEs, we propose a fast and stable training method. First, we suggest straightening the sampling trajectory and quickly finetuning the multi-step model into a one-step model by replacing the distillation loss function with the proposed feature matching loss. Then, we extend the Diff-Instruct training strategy [37], using the gradient of the proposed feature matching loss to replace the gradient provided by score distillation in the latter half of the timestep. We name our proposed method SDXS to admire SDXL [2], and provide two versions: $\times 512$ 512 and $\times 1024$ 1024, developed based on SD v2.1 base and SDXL, respectively. As shown in Tables 1 and 2, SDXS demonstrates efficiency far surpassing that of the base models, even achieving image generation at 100 FPS for $\times 512$ 512 images and 30 FPS for $\times 1024$ 1024 images on the GPU. Finally, to ease the application of our optimized model to tasks involving image-conditioned generation, we have adapted the block removal distillation strategy for ControlNet [5]. Then, we extend our proposed training strategy to the training of ControlNet, relying on adding the pretrained ControlNet to the score function.
@@ 2.2 Diff-Instruct
where $\bm{x}_{0}=g_{\theta}(\bm{z})$ denotes the sample generated by the one-step generator being trained to accept a randomly initialized latent $z$, and $\bm{s}_{\phi}$ and $\bm{s}_{p_{t}}$ denote the score functions of the DM trained online on the generated data and the pretrained DM, respectively. Diff-Instruct uses this gradient directly to update the generator, and when the outputs of the two score functions agree, the marginal distribution of the one-step generator output is consistent with the marginal distribution of the pretrained DM. It’s worth noting that the core concept of Diff-Instruct closely aligns with that of VSD [14], and this concept has also been recently expanded to text-to-image generation tasks [40, 41].
@@ 3.1 Architecture Optimizations
where $\lambda_{F}$ balances the two loss terms. Different from BK-SDM, we exclude the original denoising loss. Our models are miniaturized based on SD-2.1 base and SDXL-1.0 base. For SD-2.1 base, we remove the middle stage, the last stage of the downsampling stages and the first stage of the upsampling stages, and remove the Transformer blocks of the highest resolution stages. For SDXL-1.0 base, we remove most of the Transformer blocks.
@@ 3.2 One-Step Training
A straightforward approach to training a one-step model involves initializing noises $\bm{\epsilon}$ and employing an Ordinary Differential Equation (ODE) sampler $\psi$ to sample and obtain the generated images $\bm{\hat{x}}_{0}$, thereby constructing noise-image pairs. These pairs then serve as inputs and ground truth for the student model during training. This method, however, often results in the production of low-quality images. The underlying issue is the crossings in the sampling trajectories of noise-image pairs generated using the ODE sampler from a pretrained DM, leading to an ill-posed problem. Rectified Flow [29, 30] tackles this challenge by straightening the sampling trajectories. It replaces the training objective and proposes a ’reflow’ strategy to refine the pairing, thus minimizing trajectory crossings. Instead, we note that the crossing of sampling trajectories can lead to a one noise input corresponding to multiple ground-truth images, causing the trained model generating an image that is a weighted sum of multiple feasible outputs with weights $w(y)$:
@@
Input: offline DM $\bm{s}_{p_{t}}$, online DM $\bm{s}_{\phi}$, ODE sampler $\psi$, prior distribution $p_{z}$, one-step DM $\mathbf{x}_{\theta}$, hyperparameters $\lambda_{FM}$ and $\alpha$.
while not converge do
     update ${\phi}$ using SGD with gradient
$\operatorname{Grad}(\phi)=\frac{\partial}{\partial\phi}\int_{t=0}^{T}w(t)\mathbb{E}_{\mathbf{z}\sim p_{z},\bm{x}_{0}=\mathbf{x}_{\theta}(\mathbf{z}),\atop\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\|\bm{s}_{\phi}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{x}_{0})\|_{2}^{2}\mathrm{d}t.$
update $\theta$ using SGD with the gradient
$\operatorname{Grad}(\theta)$ $=\int_{t=0}^{\alpha T}w(t)\mathbb{E}_{\bm{x}_{0}=\mathbf{x}_{\theta}(\bm{z}),\bm{x}_{t}|\bm{x}_{0}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})}\big{[}\bm{s}_{\phi}(\bm{x}_{t},t)-\bm{s}_{p_{t}}(\bm{x}_{t})\big{]}\frac{\partial\bm{x}_{t}}{\partial\theta}\mathrm{d}t$
$+\lambda_{FM}\frac{\partial\mathcal{L}_{FM}(\mathbf{x}_{\theta}(\bm{x}_{t}),\psi(\mathbf{x}_{\phi}(\bm{x}_{t})))}{\partial\theta}.$
adjust hyperparameters $\lambda_{FM}$ and $\alpha$.
end while
return $\theta,\phi$.
Algorithm 1 Segmented Score Distillation Algorithm
Segmented Score Distillation.
@@
$\lambda_{FM}$ is used to balance gradients of the two segments, and $\alpha\in[0,1]$. we intentionally set $\alpha$ close to 1 and $\lambda_{FM}$ at a high value to ensure the model’s output distribution smoothly aligns the predicted distribution by the pretrained score function. After achieving significant overlap in the probability densities, we gradually lower both $\alpha$ and $\lambda_{FM}$. Figure 3 visually depicts our training strategy, where the offline DM represents the U-Net of a pretrained DM, and the online DM is initialized from the offline DM and finetuned on the generated images through Eq. (1). In practice, the online DM and student DM are trained alternately as in Algorithm 1.
LoRA.
Once the one-step DM is trained, it can be finetuned like other DMs to adjust the style of the generated images. We utilize LoRA [34] in conjunction with the proposed Segmented Score Distillation to finetune an one-step DM, as illustrated in Figure 4. Specifically, we insert the pretrained LoRA into the offline DM, and if it is also compatible with the Teacher DM, it is inserted there as well. It is important to note that we do not insert LoRA into the online DM, as it corresponds to the output distribution of the one-step DM. Then, we use the same training procedure as we do for one-step training, but skip the feature matching warmup, since LoRA finetune is much more stable than full finetune. Besides, when the Teacher DM cannot incorporate the pretrained LoRA, we use a reduced $\lambda_{FM}$. In this way, the pretrained LoRA can be distilled into the SDXS’s LoRA.
Figure 4: The proposed LoRA training strategy based on feature matching and score distillation. The dashed lines indicate the gradient backpropagation.
@@
Our approach can also be adapted for training ControlNet, enabling the tiny one-step model to incorporate image conditions into its image generation process, as depicted in Figure 5. Compared to the base model for text-to-image generation, the model trained here is the distilled ControlNet that accompanies the tiny U-Net mentioned earlier, and the parameters of the U-Net are fixed during training. Importantly, we need to extract the control images from the images sampled by the teacher model, rather than from the dataset images, to ensure that the noise, target image, and control image form a pairing triplet. Furthermore, the original multi-step U-Net’s accompanying pretrained ControlNet is integrated with both the online U-Net and offline U-Net but does not participate in training. Similar to the text encoder, the function is confined to serving as a pretrained feature extractor. In this way, to further reduce $\mathcal{L}$, the trained ControlNet is to learn to utilize the control images extracted from the target images. At the same time, the score distillation encourages the model to match the marginal distributions, enhancing the contextual relevance of the generated images. Notably, we found that replacing a portion of the noise input to U-Net with freshly reinitialized noise can enhance control capabilities.
@@ 4 Experiment
Implementation Details. Our code is developed based on diffusers library [^1] https://github.com/huggingface/diffusers. Because we cannot access the training datasets of SD v2.1 base and SDXL, the entire training process is almost data-free, relying solely on prompts available from publicly accessible dataset [58]. When necessary, we use open-source pretrained models in conjunction with these prompts to generate corresponding images. To train our model, we configure the training mini-batch size to range from 1,024 to 2,048. To accommodate this batch size on the available hardware, we strategically implement gradient accumulation when necessary. It’s important to note that we found that the proposed training strategy results in models generating images with less texture. Therefore, after training, we utilize GAN loss in conjunction with extremely low-rank LoRA for a short period of fine-tuning. When GAN loss is needed, we use Projected GAN loss from StyleGAN-T [59], and the basic settings are consistent with ADD [32]. For the training of SDXS-1024, we use Vega [17], a compact version of SDXL, as the initialization for both online DM and offline DM to reduce training overhead.
@@
SD v1.5 [1] $\times 512$ 512 860 M DPM-Solver++ 16 276 24.28 31.84
SD v1.5-LCM [28] $\times 512$ 512 860 M LCM 4 84 34.74 30.85
SD Turbo [32] $\times 512$ 512 865 M - 1 35 26.50 33.07
Tiny SD [16] $\times 512$ 512 324 M DPM-Solver++ 16 146 31.16 31.06
SDXS-512 (ours) $\times 512$ 512 319 M - 1 9 28.21 32.81
@@
SDXL [2] $\times 1024$ 1024 2.56B Euler 32 1869 24.60 33.77
SDXL-LCM [28] $\times 1024$ 1024 2.56B LCM 4 299 28.87 32.45
SDXL Lightning [33] $\times 1024$ 1024 2.56B - 1 131 29.26 32.09
Vega [17] $\times 1024$ 1024 0.74B Euler 32 845 29.24 32.99
SDXS-1024 (ours) $\times 1024$ 1024 0.74B - 1 32 30.92 32.32
@@ 4.1 Text-to-Image Generation
We report the quantitative results, i.e., FID [61] and CLIP scores [62], on MS-COCO 2017 validation set [63] for evaluation. Due to the strong assumption of Gaussian distribution, FID is not a good indicator for measuring image quality [2], as it is significantly affected by the diversity of generated samples, but we still report it as prior works do. Table 3 shows the performance comparison on MS-COCO 2017 5K subset and Figure 7 shows some examples. Despite a noticeable downsizing in both the sizes of the models and the number of sampling steps required, the prompt-following capability of SDXS-512 remains superior to that of SD v1.5. Moreover, when compared to Tiny SD, another model designed for efficiency, the superiority of SDXS-512 becomes even more pronounced. This observation is consistently validated in the performance of SDXS-1024 as well. Besides, we demonstrate the sample of training LoRA using the proposed method, as shown in Figure 9. It is evident that the style of the generated images by the model can be effectively transferred to match the style of the offline DM into which the style-oriented LoRA has been integrated, while generally maintaining the consistency of the scene’s layout.
Figure 9: Two samples of two different styles of LoRA with top and bottom images generated using the same prompts.
@@ 5 Conclusion
This paper explores the distillation of large-scale diffusion-based text-to-image generation models into efficient versions that enable real-time inference on GPUs. Initially, we employ knowledge distillation to compress both the U-Net architecture and the image decoder. Subsequently, we introduce a novel training strategy that leverages feature matching and score distillation to reduce the sampling process to one step. This approach allows for the real-time generation of $\times 1024$ 1024 images on a single GPU, maintaining quality comparable to original models. Moreover, the training methodology we propose can also adapt to tasks involving image-conditioned generation, eschewing the direct adaptation of the pretrained ControlNet. We believe that the deployment of efficient image-conditioned generation on edge devices represents a promising avenue for future research, with plans to explore additional applications such as inpainting and super-resolution.